OpenStudy (anonymous):

Find the doubling time of a quantity that is increasing by 9.8% per year.

OpenStudy (haseeb96):

lets suppose time = x then double time = 2x okay ???

OpenStudy (anonymous):

t = log 2/log 1.08 = 9.006468 years

OpenStudy (anonymous):

Why log 1.08?

OpenStudy (welshfella):

yea where did you get that from?

OpenStudy (anonymous):

Somebody help, I'm confused :(

OpenStudy (anonymous):

i gave you the answer i cant give you the show work to

OpenStudy (welshfella):

oops guys I must apologise I had it the wrong way round! you divide the rate into 72 so it is 72 / 9.8 = 7.35 years

OpenStudy (anonymous):

That's wrong @welshfella

OpenStudy (anonymous):

@kinggg How did you get log 1.08? Where did the number 1.08 come from?

OpenStudy (welshfella):

how do you know its wrong?

OpenStudy (anonymous):

I entered it and it said answer is NOT CORRECT

OpenStudy (anonymous):

@kinggg 9.006468 years is incorrect too

OpenStudy (welshfella):

i think the log formula is ln2 / ln(1.098)

OpenStudy (welshfella):

that gives 7.41 years

OpenStudy (anonymous):

7.41 years is correct!

OpenStudy (welshfella):

Lets do it the long way teh suppose the principle is $100 we have double amount 200 = 100(1 + 0.098)^n (1.098)^n = 2 take logs:- n ln 1.098 = ln n = ln 2 / ln1.098) = 7.41

OpenStudy (anonymous):

are you doing Exponential growth, doubling time

OpenStudy (welshfella):

rounding that up its 8 years

OpenStudy (anwaarullah):

8 years ,, let 100 be the investment so 100*1.098=,,,,,

OpenStudy (welshfella):

yea

OpenStudy (welshfella):

i guess that the interest is paid at the end of the year so that would make it 8.

OpenStudy (anonymous):

Thanks all!

OpenStudy (welshfella):

yw